Sarah had been contributing $300 pre-tax per month to a retirement account that pays 2.16% interest compounded monthly. After 10 years, she needs to withdraw 25% of the money from her account. If the early withdrawal penalty is 10% of the amount withdrawn, how much will she have to pay?

a. $1,003.59 b. $40,143.57 c. $4,014.36 d. $10,035.89

Sergio plans to retire in 15 years. He would like to have $250,000 in his retirement account. If he invests in a plan that pays 4.69% interest compounded monthly, how much should he contribute monthly?

a. $138.89 b. $277.88 c. $959.77 d. $560.23


1
Expert's answer
2021-11-18T17:44:44-0500

a)"Total Payments = A\times nper"


Total payments="\$300\times12\times10=\$36,000"


Future value, F"= A\times\frac{(1+rate)^{nper }- 1}{rate }"


"rate = (1+\frac{r}{n})^{\frac{n}{p}}-1"


"nper = p \times t"


Where;

  • r = nominal annual interest rate
  • n = number of compounding periods per year
  • p = number of payment periods per year
  • rate = rate per payment period
  • nper = total number of payment periods
  • A = an amount added to the principal at the end of each payment period


Future value, FV"= 300\times\frac{(1.0018)^{120 }- 1}{0.0018}=40143.57"


Withdrawing "25\%" and incurring a "10\%" withdrawal fee, she will pay:


"0.25\times0.1\times40143.5=\$1003.59"


Sarah will have to pay $1003.59 to withdraw "25\%" of the money from her account.


b)

Future value, FV"= A\times\frac{(1+rate)^{nper }- 1}{rate }"


"250000= A\times\frac{(1+0.0039)^{180 }- 1}{0.0039}"


"250000= 260.47A"


"\therefore A=\$959.77"


Sergio will contribute $959.77 monthly in order to have "\$250,000" in his retirement account after 15 years.